# Consumption Smoothing in Metropolis: Evidence from the Working-class Households in Prewar Tokyo

### 26 Jan 2024 v3

@@ B.6.2 Ancillary Wage
Tsukishima Tokyo City
(1) Number (2) Average (3) Average (4) Average
of factories number of share of share of
Factory size factory workers ancillary wage (%) ancillary wage (%)
Total 47 124 12 9
10–99 workers 39 32 9 8
100–199 workers 5 143 18 17
200+ workers 3 1284 38 31
Panel C: Distribution of the Factory Size
Number of Machinery Factories (4) THBS heads
Factory size (1) THBS heads (2) Tokyo City (3) Tsukishima Ancillary wage (%)
Total 29 846 39 35
10–99 workers 6(21) 760 (90) 39(83) 10
100–199 workers 8(28) 52(6) 5(11) 19
200+ workers 15(52) 34(4) 3(6) 59
Notes: Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21). Panel B: Columns 1 and 2 summarize the number of factories and average number of workers surveyed in 1920 (Handbook of Factories 1921). Columns 3 and 4 show the average share of ancillary monthly wage predicted in Figure B.9 in Tsukishima and Tokyo City, respectively. Panel C: Column 1 summarizes the number of THBS heads in each factory size bin. One head whose factory size is unknown is not included. Two heads whose factory names are unavailable are not included. Three heads with own businesses are classified into the 10–99 workers bin. Two heads working in the other sector are excluded. Columns 2 and 3 show the machinery factories measured in the Handbook of Factories (1921 edition). Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). Data on the number of heads are from the THBS sample.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
where $m\in\{\text{December}1918,\text{January}1919,...,\text{December}1919\}$ is the year-month cell and $n_{m}$ is the frequency weight for year-month cell $m$.
Substitute equation 1011, and 12 into equation 9 to yield the final specification:
@@
Panel A in Table B.4 summarizes all the components in equation 13. In column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is roughly $52.9$ yen, which is approximately $3.5$ yen smaller than the average monthly earnings in the THBS heads ($56.4$ yen, see Panel C of Table 2) as I expected in Online Appendix B.6.2. However, the deviation from the population mean (i.e., $3.5$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $3.5$ yen difference is less than two days’ wages. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Another reason for supporting the representativeness is the systematic downward bias in estimating the average daily wage explained in Online Appendix B.6.1. When I calculated the average daily wage of the male factory workers in the skilled age range, I used the wage statistics for the machinery sector in Tokyo City, given the limited availability of the wage for age bin statistics. Since the average size of factories in Tsukishima was slightly larger than in the entire city (Panel C in Table B.3), using the wage statistics for the whole of the city systematically underestimated the average daily wage in Tsukishima. Thus, $52.9$ yen is a conservative estimate for the average monthly wage for the skilled workers in the machinery sector in Tsukishima.
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $51.4$ yen, which is smaller than that in column 1, reflecting the difference in the weights for the ancillary wages. When I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in the deviation from the THBS heads’ earnings (i.e., $5$ yen).
B.7 Measurement Errors in the Raw and Shifted Semi-monthly Dataset
Figure B.10: Comparisons of Average Monthly Incomes
@@ B.7 Measurement Errors in the Raw and Shifted Semi-monthly Dataset
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from raw, shifted, and adjusted panel datasets, respectively. Raw series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half”. As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can deal with the measurement errors from the timing of payday but can not adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series can deal with measurement errors and take the timing of the household expenses into account. Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from raw, shifted, and adjusted panel datasets, respectively. Source: Created by the author using the THBS sample.
Figure B.10: Comparisons of Average Monthly Incomes
@@
Raw series defines the days from the 1st to the 15th of each month as the first half and the days from the 16th to the end of each month as the second half. This raw series shall contain substantial measurement errors due to the shifts of paydays explained in Section 3. The time-series plots of average income in Figure B.10a have substantially large variance and fluctuate over time. This is because some semi-monthly income is assigned to the wrong semi-month cells. In these wrong cells, the measured incomes shall take larger values than the actual ones. Moreover, the time-series plots of average expenditure do not chase those of the income. For example, the average expenditure in the second half of June shows a large difference from the average income in the same cell. Consequently, the average net income (i.e., income minus expenditure) in each semi-month is relatively flat and takes a few odd cells with large positive values (Figure B.10b).
@@
The shifted series defines the days from the 2nd to the 16th of each month as the first half and the days from the 17th to the 1st of the following month as the second half. Although the shifted series can deal with the measurement errors from the payday shifts, it can not adjust the difference in the payday timing and expenses. Figure B.10c indeed gets a much smaller variance in the average incomes over time because all the paydays are now fixed to assign the suitable semi-month cells. However, the expenditure series still does not chase the income series well because the payday is placed on the latter edge in each semi-month cell. In other words, the average expenditure in each semi-month reflects a large part of the consumption based on the income of the previous semi-month cell. Since the average incomes take a better series than Figure B.10a, the average net income now seems to show an increasing trend (Figure B.10d). However, it still has a few cells with large negative and large values due to the deviations of expenditure from income.
@@
The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half”. By shifting the entire semi-month cells, this adjusted series can deal with the measurement errors from the payday shifts and fix the miss-matching in the timing of income and expenditure. In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption given the latest income. Figure B.10e shows similar time-series plots of average income in Figure B.10c, in which has a corrected semi-monthly average income series. Notably, the expenditure series are now chasing the income series in Figure B.10e because each semi-month cell’s expenditure is based on the income in the same cell. Accordingly, Figure B.10f now shows a smoother series in the net income. It shows a clear increasing trend in the net income and has a smaller variance over time.
@@
These illustrations support the evidence that the adjusted semi-monthly dataset can offer cleaner within variations in income and expenditure because it can fix the measurement errors raised from the shifts of paydays and can reflect the households’ consumption behaviors based on their latest incomes measured in the same semi-month cell.
@@ B.8 Distribution of the Idiosyncratic Shocks
Notes: This figure shows the distribution of idiosyncratic shocks in the semi-monthly THBS panel dataset. The idiosyncratic shocks are estimated as the residuals from the regression of the first difference in income on the first difference in aggregate income. The red line indicates the mean of the idiosyncratic shocks ($-0.369$). Source: Created by the author using the THBS sample.
@@
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset. [^87] I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the denser panels include a set of consecutive household-date cells with no income. These cells can not be used to predict the idiosyncratic shocks as the residuals economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{j}$ be the number of households in semi-month $j$ in the year 1919. For household $i$ in semi-month $j$ cell, the residual is then defined as follows:
@@
It is clear that the residuals satisfy the zero-average property (Figure B.11). However, the residuals have some variations, indicating that the THBS households had experienced both favorable and adverse idiosyncratic shocks during the survey periods.
@@ B.9 Unit Root and Autocorrelations
Table B.5 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and adjusted monthly (Column 2). Both columns show that the null hypothesis of unit roots in all the panels is rejected at the conventional level. The results of the unit root tests for consumption subcategories are not reported as those are materially similar results. Since I used the cluster-robust variance-covariance estimator, any arbitrary autocorrelations within unit are allowed in the regressions (Arellano 1987). Despite this, I confirmed that the null hypothesis of no serial correlation is not rejected at the conventional level in Wooldridge’s autocorrelation test for almost of the consumption and net income variables (Wooldridge 2002).
@@ B.10 Risk Preferences
The potential influences of heterogeneous risk preference on standard risk-sharing regressions are discussed in a few works. In my empirical setting, the heterogeneity in the risk preference is assumed to be removed by using the household fixed-effect. When the risk preferences are difference over households, say $\sigma=\sigma_{i}$, equation 24 becomes:
@@
Schulhofer-Wohl (2011) shows that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised under the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock can not be removed using the time-fixed effect in the risk-sharing regressions. [^88] Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (16) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (17) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 17. As explained, this is a different setting from the one used in the current paper. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that can not be applicable to my THBS dataset. [^89] The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
@@
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks are different across high- and low-risk-tolerance workers. [^90] This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there’s a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
Since the THBS dataset does not have any experimental questions, it is impossible to apply a similar test under my empirical setting. Considering the limitation that the direct test using experimental risk-tolerance question is unavailable, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to get insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
$\log\tilde{y}_{i,t}=\xi_{1}\log c_{t}^{a}+\xi_{2}\log c_{t}^{a}\times d_{i}+ox_{i,t}+\iota_{i}+\varepsilon_{i,t},$ (18)
@@
where $\tilde{y}_{i,t}$ is head’s earnings, $c_{t}^{a}$ is aggregate consumption, $x_{i,t}$ is the household size, $\iota_{i}$ is the household fixed-effect, and $\varepsilon_{i,t}$ is a random error term. The aggregate consumption is calculated as the cross-sectional average of the total expenditure among THBS sample, given that the mean consumption expenditure among all the skilled factory workers’ households in Tsukishima is unavailable.
@@
$d_{i}$ is an indicator variable that takes one if household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I consider the case that the heads who worked in the large-scale factory had different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factory might have had a lower risk tolerance than those in the small- and medium-scale enterprises. 13 heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I assume that the heads who worked in the smithing and the non-machinery factories had different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be business owners. [^91] Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
@@
Table B.6 presents the results. In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant. Thus, the sensitivities to aggregate consumption are similar across different scales of the factories. Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance. It suggests that the heads worked in the smithing and non-machinery factories show similar responses to the aggregate shock. Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations. To deal with this issue, I exclude the semi-months in October and November 1919 in Columns (3) and (4). Clearly, the estimated coefficients on the interaction terms are largely unchanged in both columns.
@@
This finding is consistent with the fact that these heads have similar characteristics in nature because the budget survey focused on skilled factory workers in Tsukishima. Thus, it would be plausible to see that the THBS households had more or less similar attitudes toward the risks.
@@
Aggregate Consumption 0.311* 0.297* 0.378** 0.372**
[0.154] [0.174] [0.154] [0.171]
Aggregate Consumption -0.080 -0.080
$\times$ Large Scale Factory [0.400] [0.366]
Aggregate Consumption -0.101 -0.176
$\times$ Smithing/Non-machinery [0.191] [0.188]
Observations 263 263 239 239
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 18 for the semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed-effect. Columns 1–2 show the results for the entire sample period. Columns 2–3 show the results for the period between January and September 1919 to remove year-month cells with the smaller number of households (Online Appendix Figure B.4).

### 25 Feb 2024 v6

@@ B.6.2 Ancillary Wage
(1) Tokyo City (2) Tsukishima
Number Number Average share Number Number Average share
of factories of factory of ancillary of factories of factory of ancillary
Factory size workers wage (%) workers wage (%)
10–99 workers 760 18,760 8 39 1,252 9
100–199 workers 52 7,155 17 5 713 18
200+ workers 34 21,764 31 3 3,851 38
Total 846 47,679 9 47 5,816 12
Panel C: Comparisons of the Distributions of Factory Size
Machinery Factories
(1) Tokyo City (2) Tsukishima (3) THBS
Number % share Number % share Number % share
of factory of factory of THBS
Factory size workers workers heads
10–99 workers 18,760 39 1,252 22 6 21
100–199 workers 7,155 15 713 12 7 25
200+ workers 21,764 46 3,851 66 15 54
Total 47,679 100 5,816 100 28 100
Notes: Panel A: This panel shows the size of factories and the share of fixed and ancillary monthly wages in six factories surveyed in Kitazawa (1924). The number of factory workers surveyed in factories A–F are 14, 23, 9, 11, 11, and 16, respectively. The share of these surveyed skilled workers to all skilled workers is 20.3 in Factory A, 50.0 in Factory B, 5.3 in Factory C, 2.3 in Factory D, and 4.2 in Factory F, respectively. This figure for factory E is unavailable because the number of skilled workers is unknown (Kitazawa 1924, p. 21). Panel B: Subcolumns 1–2 and 4–5 summarize the number of factories and total number of workers in Tokyo City and Tsukishima, respectively (Handbook of Factories 1921). Subcolumns 3 and 6 show the average share of ancillary monthly wage predicted in Figure B.9 for Tokyo City and Tsukishima, respectively. Panel C: Columns 1–3 summarize the number of factory workers and its share in percentage points in Tokyo City, Tsukishima, THBS sample, respectively. The statistics in Columns 1 and 2 (same figures used in Panel B) are from the Handbook of Factories (1921). In Column 3, the THBS heads who worked in the machinery factories are considered. There are 28 heads whose factory names and size are available. For a blacksmith included in these heads, I assume the number of workers to be 10, based on the report (Department of Health, Ministry of the Interior 1923a, p. 412). Sources: Panel A: Kitazawa (1924, pp. 20; 58–62). Panel B: The number of factory workers is taken from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). The average shares of ancillary wage are calculated using Kitazawa (1924, pp. 20; 58–62). See also Figure B.9. Panel C: Data on the number of factory workers and factories are from the Ministry of Agriculture and Commerce (1921, pp. 613-629; 674). Data on the number of heads are from the THBS sample.
@@ B.6.3 Estimating Monthly Income of Adult Male Factory Workers
where $m\in\{\text{December}~{}1918,\text{January}~{}1919,...,\text{December}~{}1919\}$ is the year-month cell and $n_{m}$ is the frequency weight for year-month cell $m$.
Substitute equation 1112, and 13 into equation 10 to yield the final specification:
@@
Panel A in Table B.4 summarizes all the components in equation 14. In column 1 of Panel B in Table B.4, I calculate the average monthly wage under the share of the ancillary wage for the factories in Tsukishima ($\vartheta$). The estimate is $55$ yen, which is approximately $2$ yen higher than the average monthly earnings in the THBS heads ($53$ yen, see Panel C of Table 2) as I expected in Online Appendix B.6.2. Therefore, the deviation from the population mean ($55$ yen) is not too large to infer the mean household behaviors of the skilled workers in the machinery factories in Tsukishima. Given that the average daily wage of the male factory workers in the skilled age range is $2.5$ yen (Figure B.8), $2$ yen difference is less than one day’s earnings. This is consistent with the findings in my descriptive analyses that show the similarities in the distribution of family structure, average household size, and share of ancillary wage between the THBS households and Tsukishima households (Section 3; Online Appendix B.6.2).
Column 2 uses the share of the ancillary wage for the entire Tokyo City ($\vartheta=0.09$) to calculate the average monthly wage. The estimate is $53.6$ yen, which is slightly smaller than that in Column 1, reflecting the difference in the weights for the ancillary wages. When I see this figure as an unbiased estimate for the entire city, the difference in the average factory size between the entire Tokyo City and Tsukishima may be reflected in this deviation (Panel B in Table B.3).
B.7 Measurement Errors in the Unadjusted Panel Dataset
Figure B.10: Comparisons of Average Semi-monthly Income and Expenditure
@@ B.7 Measurement Errors in the Unadjusted Panel Dataset
Note: Figures B.10aB.10cB.10e show the average semi-monthly income and expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Unadjusted series defines the days from the 1st to the 15th of each month as the first half, whereas the days from the 16th to the end of each month as the second half. The shifted series defines the days from the 2nd to the 16th of each month as the first half, whereas the days from the 17th to the 1st of the following month are the second half. Adjusted series is the baseline definition in the main text, which uses the days from 12th to 26th in each month as the “first-half” and those from 27th to 11th in the next month as the “second-half”. As explained in Section 3, the raw series contains considerable measurement errors due to the timing of payday. The shifted series can deal with the measurement errors from the timing of payday but can not adjust the difference in the payday timing and expenses. In other words, the shifted series ignores the households’ decision on the consumption schedule in each semi-month. The adjusted series can deal with measurement errors and take the timing of the household expenses into account. Figures B.10bB.10dB.10f illustrate the average semi-monthly income minus expenditure calculated from unadjusted, shifted, and adjusted panel datasets, respectively. Source: Created by the author using the THBS sample.
Figure B.10: Comparisons of Average Semi-monthly Income and Expenditure
@@
The unadjusted (calender-month) series defines the days from the 1st to the 15th of each month as the first half and the days from the 16th to the end of each month as the second half. This series shall contain substantial measurement errors due to the mismatches in the income and expenditure data explained in Section 3.3. The time-series plots of average income in Figure B.10a have substantially large variances and fluctuate over time. This is because some semi-monthly income is assigned to the wrong semi-month cells. In these wrong cells, the measured incomes shall take unnaturally greater values than usual. Moreover, the time-series plots of average expenditure do not chase those of the income. For example, the average expenditure in the second half of June greatly differs from the average income in the same cell. Consequently, the average net income (i.e., income minus expenditure) in each semi-month is relatively flat and takes a few odd cells with large positive values (Figure B.10b).
@@
The shifted series defines the days from the 2nd to the 16th of each month as the first half and the days from the 17th to the 1st of the following month as the second half. Although the shifted series can partly deal with the measurement errors in income, it does not adjust the difference in the payday timing and expenses. Figure B.10c indeed gets a much smaller variance in the average incomes over time because all the paydays are now fixed to assign the suitable semi-month cells. However, the expenditure series still does not chase the income series well because the payday is placed on the latter edge in each semi-month cell. In other words, the average expenditure in each semi-month reflects a large part of the consumption based on the income of the previous semi-month cell. Since the average incomes take a better series than Figure B.10a, the average net income offers better time-series plots (Figure B.10d). However, it still has a few cells with large negative and large values due to the deviations of expenditure from income.
@@
The adjusted series uses the days from the 12th to 26th of each month as the “first half” and those from the 27th to 11th of the following month as the “second half”. By shifting the entire semi-month cells, this adjusted series can deal with the measurement errors from the payday shifts and fix the miss-matching in the timing of income and expenditure. In other words, each payday is assigned to a suitable semi-month cell, and each semi-month cell reflects the household’s decision on consumption given the latest income. Figure B.10e shows similar time-series plots of average income in Figure B.10c, in which has a corrected semi-monthly average income series only. Notably, the expenditure series are now chasing the income series in Figure B.10e because each semi-month cell’s expenditure is based on the income in the same cell. Accordingly, Figure B.10f now shows a smoother series in the net income. It shows a clear increasing trend in the net income and has a smaller variance over time.
@@
These illustrations support the evidence that the adjusted semi-monthly datasets can offer cleaner within variations in income and expenditure because they can fix the measurement errors and can reflect the households’ consumption behaviors based on their latest incomes measured in the same semi-month cell.
@@ B.8 Distribution of the Idiosyncratic Shocks
Notes: This figure shows the distribution of idiosyncratic shocks in the semi-monthly THBS panel dataset. The idiosyncratic shocks are estimated as the residuals from the regression of the first difference in income on the first difference in aggregate income. The red line indicates the mean of the idiosyncratic shocks ($-0.235$). Source: Created by the author using the THBS sample.
@@
Figure B.11 illustrates the distribution of the idiosyncratic shocks predicted using the semi-monthly panel dataset. [^95] I used the semi-monthly panel dataset to overview the entire distribution of the idiosyncratic shocks rather than the daily panel datasets. This is because the denser panels include a set of consecutive household-date cells with no income. These cells can not be used to predict the idiosyncratic shocks as the residuals economically meaningfully, and thus it is practically impossible to illustrate the idiosyncratic shocks based on the daily dataset. The idiosyncratic shocks are calculated as the residuals from the regression in the first difference in semi-monthly income on the first difference in aggregate semi-monthly income. Let $n_{t}$ be the number of households in semi-month $t$ in the year 1919. For household $i$ in semi-month $t$ cell, the residual is then defined as follows:
@@ B.9 Unit Root and Autocorrelations
Table B.5 presents the results of the Fisher-type panel unit root tests for the total expenditure in the semi-monthly (Column 1) and adjusted monthly (Column 2) panel datasets. Both columns show that the null hypothesis of unit roots in all the panels is rejected at the conventional level. The results of the unit root tests for consumption subcategories and variables used to test the risk-coping mechanisms are not reported in this table because those are materially similar results. In Section 4, I used the cluster-robust variance-covariance estimator in the regressions to allow arbitrary autocorrelations (Arellano 1987). Despite this, I confirmed that the null hypothesis of no serial correlation is not rejected at the conventional level in Wooldridge’s autocorrelation test for almost of the consumption and net income variables (Wooldridge 2002).
@@ B.10 Risk Preferences
The potential influences of heterogeneous risk preference on standard risk-sharing regressions are discussed in a few works (Schulhofer-Wohl 2011; Mozzoco and Saini 2012). In my empirical setting, the heterogeneity in the risk preference is assumed to be removed by using the household fixed-effect. When the risk preferences are difference over households, say $\sigma=\sigma_{i}$, equation 26 becomes:
@@
Schulhofer-Wohl (2011) shows that the heterogeneous risk preference may lead to bias in the estimation of the income elasticity under the standard risk-sharing regressions and proposes a parametric test for that heterogeneity. This test is valid under a particular utility function. Specifically, the bias could be raised under the case that the coefficient of relative risk aversion interacts with the aggregate shock in the first-order condition for consumption because the interaction means that the aggregate shock can not be removed using the time-fixed effect in the risk-sharing regressions. [^96] Schulhofer-Wohl (2011, pp. 928–929) considers the utility maximization problem under the following weighted sum of discounted expected utilities: $\sum_{i}\dot{\alpha}_{i}E\left[\sum_{t}\dot{\beta}^{t}\frac{[c_{i,t}(s_{t})]^{1-\dot{\gamma}_{i}}}{1-\dot{\gamma}_{i}}\right],$ (17) where $\dot{\alpha}_{i}$ is Parate weight and $\dot{\gamma}_{i}$ is the coefficient of relative risk aversion for household $i$. The log-transformed first-order condition becomes: $\text{log}c_{i,t}=\frac{\text{log}\dot{\alpha}_{i}}{\dot{\gamma}_{i}}-\frac{\text{log}\dot{\lambda}_{t}}{\dot{\gamma}_{i}},$ (18) where $\dot{\lambda}_{t}$ indicates the Lagrangian multiplier. This result implies that the effects of the aggregate shocks ($\dot{\lambda}_{t}$) depend on the household’s attitudes toward risk and that the first-differencing does not remove the second term in equation 18. As explained, this setting differs from the one used in this current study. Mozzoco and Saini (2012) consider a more general case for testing the homogeneity of risk preferences. Their test requires a panel dataset with a very long time dimension that can not be applicable to my THBS dataset. [^97] The data used in their study covers the household-level monthly panel dataset surveyed in rural India between 1975 and 1985. The households with fewer than 80 data points are not included in the homogeneity tests (Mozzoco and Saini 2012, p. 453). See also Shrinivas and Fafchamps (2018) and Mozzoco and Saini (2018) for the subsequent discussions.
@@
An alternative way to test potential heterogeneity in the risk preference is to use direct measures on the risk preferences of workers. Schulhofer-Wohl (2011) uses the results of experimental risk-tolerance questions on the workers’ risk preferences in the Health and Retirement Study to test whether the correlations between earnings and aggregate shocks differ across high- and low-risk-tolerance workers. [^98] This question is asking if the respondents would be willing to take a risk to get a new job with a higher income, even if there’s a 50-50 chance of a reduction in income by a third (Schulhofer-Wohl 2011, P. 931). They found that the high-risk-tolerance workers have more procyclical incomes than the low-risk-tolerance workers.
@@
Since the THBS has no experimental questions, it is impossible to apply the same test. Given this limitation that the direct test using experimental risk-tolerance question is unavailable, I compare the correlations between the heads’ income and aggregate consumption over sub-groups to get insights into the potential difference in the heads’ attitudes toward risks. For household $i$ and semi-month $t$, the specification is characterized as follows:
$\log\tilde{y}_{i,t}=\xi_{1}\log c_{t}^{a}+\xi_{2}d_{i}\log c_{t}^{a}+ox_{i,t}+\iota_{i}+\varepsilon_{i,t},$ (19)
@@
where $\tilde{y}_{i,t}$ is head’s earnings, $c_{t}^{a}$ is aggregate consumption, $x_{i,t}$ is the household size, $\iota_{i}$ is the household fixed-effect, and $\varepsilon_{i,t}$ is a random error term. The aggregate consumption is calculated as the cross-sectional average of the total expenditure among the THBS sample, given that the population mean of consumption expenditures among all the skilled factory workers’ households in Tsukishima is unavailable.
@@
$d_{i}$ is an indicator variable that takes one if household $i$ belongs to a specific group that might have had different preferences on their jobs than the other heads in the sample. Firstly, I consider the case that the heads who worked in the large-scale factory had different risk preferences. Although there are no specific anecdotes behind this, the workers in the large-scale factory might have had a lower risk tolerance than those in the small- and medium-scale enterprises. 13 heads who worked in the Ishikawajima shipyard and governmental factory are classified into this group. Secondly, I assume that the heads who worked in the smithing and the non-machinery factories had different risk preferences. As described in Online Appendix A.5, skilled workers in the smithing sector could be small business owners. [^99] Workers in the canning factories might have had a different preference on the risk preferences (Online Appendix A.6). However, my THBS sample does not include the workers in canning factories. The workers in non-machinery factories might also have different job preferences than the machinery factory workers. This group has only three heads: one worked in a smithing factory, and two worked in food and paper-making factories.
@@
Table B.6 presents the results. In Column (1), the estimated coefficient on the interaction term is close to zero and statistically insignificant. Thus, the sensitivities to aggregate consumption are similar across different scales of the factories. Column (2) further shows that the estimated coefficient on the interaction term is not statistically significant under an alternative proxy of risk tolerance. It suggests that the heads worked in the smithing and non-machinery factories show similar responses to the aggregate shock. Since the aggregate consumption is the cross-sectional sample average, the aggregate variable may contain noise in the period with the small number of cross-sectional observations. To deal with this potential issue, I exclude October and November 1919 in Columns (3) and (4). As shown, the estimated coefficients on the interaction terms are largely unchanged in both columns.
@@
This finding is consistent with the fact that these heads have similar characteristics in nature because the THBS focused on skilled factory workers in a representative manufacturing area, Tsukishima. Therefore, it is plausible to see that the THBS households had more or less similar attitudes toward the risks.
@@
Aggregate Consumption 0.277* 0.273 0.361** 0.362*
[0.158] [0.198] [0.164] [0.198]
Aggregate Consumption -0.041 -0.061
$\times$ Large Scale Factory [0.453] [0.415]
Aggregate Consumption -0.089 -0.178
$\times$ Smithing/Non-machinery [0.215] [0.215]
Observations 261 261 239 239
***, **, and * denote statistical significance at the 1%, 5%, and 10% levels, respectively. Standard errors in brackets are clustered at the household level. Notes: This table shows the results of equation 19 for the semi-monthly panel dataset. ‘Aggregate Consumption’ indicates the cross-sectional sample average of the total expenditures. ‘Large Scale Factory’ indicates an indicator variable that takes one if the head worked in the Ishikawajima Shipyard or governmental factory. ‘Smithing/Non-machinery’ indicates an indicator variable that takes one if the head worked in the smithing, food, or paper-making factories. All the regressions include the family size variable and household fixed-effect. Columns 1–2 show the results for the entire sample period. Columns 2–3 show the results for the period between January and September 1919 to remove year-month cells with the smaller (cross-sectional) number of households (Online Appendix Figure B.4).